### v9

# Accelerating the Exploration of Thermal Materials via a Synergistic Strategy of DFT and Interpretable Deep Learning

Abstract
Lattice thermal conductivity (LTC) is a critical parameter for thermal transport properties, playing a pivotal role in advancing thermoelectric materials and thermal management technologies. Traditional computational methods, such as Density Functional Theory (DFT) and Molecular Dynamics (MD), are resource-intensive, limiting their applicability for high-throughput LTC prediction. While AI-driven approaches have made significant strides in material science, the trade-off between accuracy and interpretability remains a major bottleneck. In this study, we introduce an interpretable deep learning framework that enables rapid and accurate LTC prediction, effectively bridging the gap between interpretability and precision. Leveraging this framework, we identify and validate four promising thermal conductors/insulators using DFT and MD. Moreover, by combining sensitivity analysis with DFT calculations, we uncover novel insights into phonon thermal transport mechanisms, providing a deeper understanding of the underlying physics. This work not only accelerates the discovery of thermal materials but also sets a new benchmark for interpretable AI in material science.
I Introduction
Lattice thermal conductivity (LTC, $\kappa_{\rm L}$) is a critical physical parameter that quantifies a material’s ability to transfer heat through lattice vibrations. It has diverse applications in thermal management [1], energy conversion [2], and thermoelectric materials [3]. High-$\kappa_{\rm L}$ materials are used in electronic devices to efficiently dissipate heat and prevent overheating [4]. In contrast, low-$\kappa_{\rm L}$ materials exhibit excellent performance in thermoelectric conversion, making them ideal for developing efficient thermoelectric generators (TEGs) [5] and coolers [6]. $\kappa_{\rm L}$ is crucial for material design and optimization, yet acquiring it for specific materials is challenging.
Traditional experimental methods for measuring $\kappa_{\rm L}$, such as the laser flash method [7] and thermal conductivity probes [8], are often inefficient. In theoretical calculations, solving the Boltzmann Transport Equation (BTE) based on Density Functional Theory (DFT) is regarded as the most reliable method for determining $\kappa_{\rm L}$ [9, 10]. Molecular Dynamics (MD) offers an alternative [11]; however, the former is limited by its substantial computational resource requirements, while the latter’s accuracy depends on the choice of interatomic potentials [12]. In recent years, machine learning (ML) has emerged as a powerful and efficient data mining tool, gaining widespread application in materials science [13]. The intersection of materials science and artificial intelligence is commonly referred to as “Materials Informatics” [14] or the “Materials Genome” [15]. Early ML-based predictions of material properties primarily aimed at achieving high prediction accuracy [16, 17, 18]. Nevertheless, such efforts, relying on “black-box” models, offered limited support for advancing theoretical research in materials science. The focus has shifted towards model interpretability, prompting greater adoption of “white-box” models that contribute more substantially to theoretical advancements [19]. Research in materials informatics related to thermal conductivity has followed this trend. Efforts to predict $\kappa_{\rm L}$ using black-box models such as Random Forest (RF) [20], Gaussian Process Regression (GPR) [21], and eXtreme Gradient Boosting (XGBoost) [22] have yielded reliable accuracy, whereas, the interpretability of these models is often hindered by their complexity. Alternatively, Genetic Programming-based Symbolic Regression (GPSR) [23] algorithms offer better interpretability, but their simplicity comes at the cost of reduced accuracy. The prevailing view is that the behavior and processes of complex models are difficult to understand and interpret, while simple models often lack strong fitting capabilities [24, 25, 19]. This presents a challenging trade-off, but the situation is gradually improving. The Sure Independence Screening and Sparsifying Operator (SISSO) [26], based on compressed sensing and symbolic regression, has not only surpassed the accuracy of the Slack semi-empirical model [27] but also narrowed the gap with black-box models like Kernel Ridge Regression (KRR) and GPR, and has been used to quantify feature sensitivity and identify key physical parameters influencing $\kappa_{\rm L}$ [28]. Despite its potential, SISSO’s applicability is constrained by its high computational demands. Enumerating combinations of features and operators to construct descriptors poses an NP-hard problem [26], with resource requirements increasing for high-dimensional input features [29].
Figure 1: A schematic illustration of the DL framework for $\kappa_{\rm L}$ modeling and high-throughput prediction. This framework facilitates accurate and interpretable predictions of $\kappa_{\rm L}$ for inorganic crystalline materials.
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving ordinary differential equations [31], time series classification [32], and fluid modeling [33]. In some cases, smaller KAN models outperform MLPs [RN165, 30]. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [RN165, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
Figure 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [34] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Furthermore, validated through DFT and MD calculations, we successfully identified three thermal conductors and one thermal insulator using this framework. Finally, taking two of these materials as case studies, we leverage DFT to provide in-depth insights into their respective phonon thermal transport mechanisms.
II Results
II.1 Establishment and evaluation of $\kappa_{\rm L}$ models
In selecting black-box models, we consider XGBoost and MLP as our primary options. XGBoost, a representative of ensemble learning, has demonstrated notable performance in various applications, including electrocaloric temperature change prediction in ceramics [17] and materials mechanical property prediction [35]. XGBoost is an enhancement of gradient boosting decision trees (GBDT). Compared to traditional GBDT, XGBoost introduces several innovations, including regularization (to improve generalization), a two-step gradient approximation for the objective function (to enhance computational efficiency), column subsampling (to reduce noise and boost generalization), and handling missing values (using a default direction for tree nodes, making it suitable for sparse datasets) [36]. MLP is a type of feedforward neural network consisting of an input layer, one or more hidden layers, and an output layer. Each neuron in a given layer is fully connected to all neurons in the preceding layer through weight matrices, allowing the network to perform linear combinations and nonlinear mappings of high-dimensional features. Neurons in each hidden layer typically use nonlinear activation functions such as ReLU, which endow the MLP with considerable expressive power, enabling it to approximate arbitrarily complex nonlinear functions. The training process of an MLP is carried out using the backpropagation algorithm, which optimizes the network’s weights through gradient descent to minimize a loss function. While MLPs exhibit substantial performance advantages in handling complex data patterns and feature learning tasks, their computational complexity and sensitivity to hyperparameter choices can make the training process time-consuming [37].
Table 1: The ranges for hyperparameter optimization.
Model Hyperparameter Range/Value
MLP hidden_sizes [64, 64, 64] (fixed)
learning_rate [0.001, 0.1]
batch_size [64, 256]
num_epochs [100, 600]
XGB n_estimators [50, 500]
learning_rate [0.0001, 0.5]
max_depth [1, 6]
subsample [0.8, 1]
colsample_bytree [0.8, 1]
reg_alpha [0.8, 1]
reg_lambda [5, 50]
KAN width [5, 15]
grid [5, 15]
k [2, 10]
lamb_l1 [5, 50]
steps [20, 35]
lr [0.01, 1.0]
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch [38] and scikit-learn [39] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna [40] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal [41]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in Table S1.
In contrast to the previously discussed algorithms, SISSO imposes the highest computational demands during model training. Therefore, we followed hyperparameter configurations recommended in established methodologies from related work [28]. To enhance interpretability, we maintained strict dimensional consistency across all features in SISSO, effectively eliminating any potential for invalid operations.
Figure 2: Model performance on primary dataset. (a) The blue circular markers represent samples from the training set, while the red square scatters are samples from the test set. (b) KAN model architecture based on the primary dataset.
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria [42]. As shown in Figure 2(a), on the primary dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low $\kappa_{\rm L}$ values (approximately $\log(\kappa_{\rm L})<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling [43] and weighting [43, 44]. However, these methods inevitably impact the model’s interpretability [45, 46]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
  • Chemical features are often non-smooth, whereas deep models tend to favor smooth solutions [47, 48];
  • Different dimensions of molecular features typically carry distinct information, yet deep models tend to integrate features across dimensions [48];
  • Tree-based models like XGBoost inherently handle redundant features, while deep models are more susceptible to interference from such features [47];
  • MLP is rotation-invariant, and any rotation-invariant learning process inherently exhibits the worst-case sample complexity [47].
In addition to accuracy, the symbolic models constructed by SISSO and KAN are also worth considering. The symbolic model provided by SISSO is as follows:
$\log{(\kappa_{\rm L})_{{\rm{SISSO}}}}=$ $-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}$
$\quad-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}$ (1)
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [49, 50], while feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$ (see Section II.2).
Table 2: Features adopted in the primary dataset.
Definition Notation Unit Range
The space group number for the relaxed structure $SG$ Unitless -
Debye temperature $\Theta$ K [24.9909, 2144.93]
Grüneisen parameter $\gamma$ Unitless [0.00210114, 3.86399]
Heat capacity per cell at constant volume (300K) ${{C}_{\text{v}}}$ $k_{\rm B}$/cell [1.11566, 275.775]
Heat capacity per cell at constant pressure (300K) ${{C}_{\text{p}}}$ $k_{\rm B}$/cell [0.0, 286.301]
Thermal expansion coefficient (300K) $\alpha$ K-1 [$-3.17\times 10^{-6}$, $5.66601\times 10^{-4}$]
[…]
Vibrational free energy per atom (300K) $F$ meV/atom [$-233.268$, 210.145]
Static bulk modulus (300K) ${{K}_{\text{s}}}$ GPa [0.0, 440.158]
Isothermal bulk modulus (300K) ${{K}_{\text{i}}}$ GPa [0.0, 437.329]
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Figure 2(b). The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq. (1). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert [51]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and $\kappa_{\rm L}$, there are a few points to consider:
  • The symbolic models constructed through machine learning require more physical features (e.g., SISSO involves 7 features, while KAN utilizes all 10 features) compared to semi-empirical models (such as Slack [27], which, when the temperature variable is fixed at 300K, requires only 5 additional features). Moreover, obtaining the corresponding features for unknown materials often involves high experimental or computational costs, making such models appear less appealing at present.
  • In materials inverse design, it is often necessary to regulate dominant features to induce changes in the target physical properties. Although the decision-making process of symbolic models is transparent, relying solely on the symbolic model itself to infer the dominance (or “contribution”) of different features is impractical.
  • Both KAN and SISSO achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Table S1), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence $\kappa_{\rm L}$. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
II.2 Deepening physical interpretability through sensitivity analysis
We identified the most critical features for $\kappa_{\rm L}$ through two steps. The first step is correlation analysis, achieved by calculating correlation coefficients. The second step is sensitivity analysis, accomplished by computing sensitivity indices.
Figure 3: Feature correlation and sensitivity analysis. (a) Heatmap of the correlation matrix. The numbers in each cell represent the Pearson correlation coefficients between the corresponding features on the horizontal and vertical axes. The color gradient visually reflects the pairwise correlation strength. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{\rm v}$) and heat capacity at constant pressure ($C_{\rm p}$), as well as isothermal bulk modulus ($K_{\rm i}$) and static (adiabatic) bulk modulus ($K_{\rm s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Figure 3(a). It is important to clarify that $K_{\rm i}$ and $K_{\rm s}$ are not entirely identical concepts [52, 53, 54]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities [52]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom ${S_{\rm{v}}}$ also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation [55]:
$F=Q-T{S_{\rm{v}}}$ (2)
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300{\rm K}$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation [56] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{\rm v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder [57], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
$H\left({\vec{\sigma},T}\right)=$ ${E_{0}}\left({\vec{\sigma}}\right)+\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T$
$+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots$ (3)
[…]
$F\approx\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T$ (4)
where $\left\langle\ln\left(\omega\right)\right\rangle\left(\sigma\right)$ is a constant, thus the vibrational free energy per atom is approximately linearly related to the Debye temperature [57]. These conclusions, combined with Eq. (4) sufficient explanation for the linear relationship between $\Theta$, $F$, and ${{S}_{\text{v}}}$.
We propose that among highly linearly correlated physical features, retaining only one is sufficient. Under ideal conditions, this approach can simplify the ML model without significantly degrading its performance. To achieve this goal, it is necessary to combine correlation analysis with sensitivity analysis. The purpose of sensitivity analysis is to determine which features have a greater impact on the target compared to others, or, more simply, which features are more “important” for the target. Currently, mainstream sensitivity analysis algorithms include SHAP [58], LIME [59], and Sobol [60]. However, these methods generally assume feature independence and the features used in our work are interdependent, their applicability is limited. Ignoring input interactions and multivariate distribution characteristics can severely skew or even invalidate any sensitivity analysis results [61]. Kucherenko et al. [62] improved the traditional Sobol index by using Copula-based sampling to separate each feature’s marginal and joint distributions, thereby constructing a dependency model. In this work, we will implement the Kucherenko indices using UQLab [63] and compare it with other indices.
As shown in Figure 3, we calculated the mean absolute values of various sensitivity indices, which provide the global values for SHAP and LIME. It is important to note that, regardless of the white-box machine learning surrogate model used, results obtained under the assumption of feature independence lack persuasiveness. In Figure 3(b), significant differences exist among the results from different methods, making it difficult to determine which is more reliable. In Figure 3(d), although different analysis methods yield similar feature sensitivities, correlated features should exhibit similar sensitivities [64]. Evidently, the conclusions from Figure 3(b) and (d) do not satisfy this premise. Therefore, conclusions drawn from sensitivity analysis methods that ignore feature interaction effects lack credibility.
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{\rm p}$ and $C_{\rm v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. (1)), specific heat capacities appear in the form $\left|{{C_{\rm v}}-{C_{\rm p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that $\log(\kappa_{\rm L})$ is entirely independent of specific heat capacities.
Figure 4: Performance comparison before and after dimensionality reduction. (a) After extracting key features and re-modeling the $\kappa_{\rm L}$, both KAN and the black-box models exhibited good robustness, while the robustness of SISSO did not meet expectations. (b) The simplified architecture of KAN model.
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $\log(\kappa_{\rm L})$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs [28, 65]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to $\log(\kappa_{\rm L})$. Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
Ultimately, we retained $C_{\rm v}$, $\alpha$, $F$, and $K_{\rm s}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of $\kappa_{\rm L}$ with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq. (12)), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
After key feature extraction, we performed symbolic regression again using KAN and SISSO. The symbolic model fitted by KAN became considerably simplified, as shown below:
$\log{\left(\kappa_{\rm L}\right)_{{\rm{KAN}}}}=$ $-1.05\sin d_{1}+0.61\cosh d_{2}-1.44$
$+1.4{{\rm{e}}^{-19.72d_{3}^{2}}}-0.48{{\rm{e}}^{-63.21d_{4}^{2}}}$ (5)
where
$d_{1}$ $=-5.66{\left({0.77-F}\right)^{2}}-0.7\tan\left({1.56{K_{\rm s}}-4.4}\right)+2.89$
$\quad+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{\rm v}}-0.11}\right)}^{2}}}}$
$d_{2}$ $=-0.76\sin\left({4.23F+4.79}\right)+1.03\tan\left({1.55\alpha-0.99}\right)$
$\quad-0.86\cosh\left({5.35{C_{\rm v}}-1.75}\right)+2.31$
$\quad+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{\rm s}}-0.63}\right)$
$d_{3}$ $={\left({0.38-{K_{\rm s}}}\right)^{3}}-0.09\sin\left({6.2{C_{\rm v}}-6.16}\right)$
$\quad-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)+0.47$
$\quad-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}}$
$d_{4}$ $=-{\left({0.37-{C_{\rm v}}}\right)^{2}}-0.13\tan\left({1.46{K_{\rm s}}-0.63}\right)-0.3$
$\quad+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}$
The analytical expression fitted by SISSO remains consistently concise:
$\log{\left(\kappa_{\rm L}\right)_{{\rm{SISSO}}}}=$ $1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}$
$-45.57\sqrt[3]{{{\alpha^{2}}{C_{\rm{v}}}{K_{\rm{s}}}}}$ (6)
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Figure 4(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between $\kappa_{\rm L}$ and the features, while also demonstrating exceptional robustness.
Figure 5: Correlation between the 4 key primary features and $\kappa_{\rm L}$. (a)-(d) Two-dimensional mappings of different features with $\log(\kappa_{\rm L})$, where the feature values have been min-max normalized. The closed solid lines represent contour lines of the sample density distribution, while the dashed lines indicate fifth-order polynomial fitting curves. (e) The mapping between $\log(\kappa_{\rm L})$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa_{\rm L})$ of each sample. The dashed line represents the mean $\log(\kappa_{\rm L})$ value of 1 for the samples.
Figure 6: Predicting key features and qualitatively assessing $\kappa_{\rm L}$ via CGCNN. (a), (b), and (c) display scatter plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets, where $R^{2}_{F}=0.940$, $R^{2}_{K_{\rm s}}=0.980$ and $R^{2}_{\alpha}=0.910$.
II.3 Screening potential thermal insulators/conductors
Unlike simple descriptors that can be directly generated by programs, such as those based on composition (e.g., Magpie [66]) or structure (e.g., Coulomb matrix [67]), the physical features involved in the aforementioned reduced dataset are relatively complex and not easily constructed physical quantities. Inspired by previous studies [68, 69], our approach is to use a two-stage prediction method. We predict primary physical features by CGCNNs, and then use the predicted primary features to perform high-throughput predictions of the $\kappa_{\rm L}$ for unlabeled materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties [34]. In our application, CGCNNs delivered fairly accurate results for the prediction of all primary features except for $C_{\text{v}}$, where the prediction accuracy was notably lower (see Supplementary Information for details).Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using primary features that can be reliably predicted qualitatively, and then calculating $C_{\text{v}}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate $\kappa_{\rm L}$.
Figure 7: $k$-means clustering and t-SNE 2D visualization. (a) Labeled training samples from AFLOW are clustered into four groups, each represented by a different color. (b) The same scatter plot as in (a), but the samples are colored according to their $\log(\kappa_{\rm L})$, with black pentagrams representing unlabeled candidates from the MP database. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the t-SNE space.
Table 3: The predicted results of $\kappa_{\rm L}$ based on KAN. The underlined materials have available DFT and MD calculation results.
MPID Formula $F_{\rm pred}$ $K_{\rm s-pred}$ $\alpha_{\rm pred}$ $C_{\rm v-DFT}$ $\kappa_{\rm L-KAN}$
mp-1009818 CN2 152.8788 333.6628 $1.71\times 10^{-5}$ 0.2343 155.2907
mp-27710 CrB4 113.7286 259.3992 $2.73\times 10^{-5}$ 12.7181 29.3775
mp-1569 Be2C 138.8222 191.1319 $2.41\times 10^{-5}$ 4.6692 80.5562
mp-1096940 CuBO2 78.15269 267.6422 $3.04\times 10^{-5}$ 8.4535 31.3254
mp-1183445 BeSiO3 85.25379 224.8873 $2.56\times 10^{-5}$ 9.5514 41.3833
mp-1184997 KRbAu2 -105.1144 14.1935 $2.36\times 10^{-4}$ 36.4862 0.2169
mp-1097263 Cs2RbNa -144.966 16.8264 $4.90\times 10^{-4}$ / /
mp-10378 Cs3Sb -143.5 11.2367 $2.42\times 10^{-4}$ / /
mp-1097633 Cs2KRb -145.791 14.89677 $5.10\times 10^{-4}$ / /
mp-635413 Cs3Bi -168.671 13.74844 $2.50\times 10^{-4}$ / /
Based on the information from Figure 3(a) and Figure 5(a)-(d), it can be concluded that $\kappa_{\rm L}$ is strongly and positively linearly correlated with both $F$ and $K_{\rm s}$. Materials with higher F and $K_{\text{s}}$ values generally exhibit higher $\kappa_{\rm L}$, a relationship that is further confirmed in Figure 5(e). Therefore, the preliminary qualitative screening of thermal insulators and conductors based on $F$ and $K_{\text{s}}$ is highly beneficial for accelerating the exploration of new materials. The predictions of $F$, $K_{\text{s}}$ and $\alpha$ using CGCNN are shown in Figure 6(a)-(c), with $R^{2}$ values reaching 0.94, 0.98 and 0.91, respectively. We extracted a total of 2,246 samples from the Materials Project (MP) [70] with the following selection criteria:
  • Excluding transition metal elements;
  • Number of atoms (nsites) $<$ 5;
  • Band gap $\in\left({0,1.5{\rm{eV}}}\right]$.
Although CGCNN is capable of accurately evaluating $F$, $K_{\rm s}$, and $\alpha$ for unlabeled materials, to the best of our knowledge, there is currently no reliable machine learning model that exhibits outstanding performance in predicting $C_{\rm v}$. When performing a quantitative estimation of $\kappa_{\rm L}$ for a large number of new materials based on Eq. (5), the $C_{\rm v}$ parameter still needs to be obtained from DFT calculations. However, this approach imposes a considerable computational cost when dealing with a large number of samples. To address this, we carry out unsupervised learning to rapidly screen the most promising thermal conductors and insulators from the MP database, enabling precise identification of candidate materials.
For this purpose, we first attempt to use $k$-means clustering [71] on samples from the AFLOW database (e.g., the labeled dataset used for model training in Section II.1) to evaluate its effectiveness. Since reliable quantitative estimation is currently only feasible for the three primary features $F$, $K_{\rm s}$, and $\alpha$, we retain only these three features for each sample during clustering. In $k$-means clustering, the optimal number of clusters is typically determined using the elbow method [72] or the silhouette score [73]. However, in our experiments, these two methods yielded inconsistent results, making it unclear which one is more convincing (see Supplementary Information). Fortunately, through manual inspection, we found that when $K=4$, the clustering results already exhibit a clear and meaningful trend.
In Figure 7(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Figure 7(a), as shown in Figure 7(b), we observe that as t-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Figure 7(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 7(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Experimental results proved that only 4 out of the 10 candidates (CN2, CrB4, Be2C and KRbAu2) effectively utilize DFT and MD to calculate $C_{\rm v}$ and $\kappa_{\rm L}$, respectively. The subsequent validation process will also focus on these 4 feasible structures.
Figure 8: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the $\kappa_{\rm L}$ to converge. The deep red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The $\kappa_{\rm L}$ calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) The blue circular scatter points and connecting lines represent the temperature dependence curve of $\kappa_{\rm L}$ for KRbAu2, while the red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model. (e) shows a histogram of the $\kappa_{\rm L}$ distribution in the training dataset.
Among the thermal conductor candidates, only Be2C belongs to the cubic crystal system. CN2 exhibits high thermal conductivity along the $z$-direction while being nearly insulating in the $x$- and $y$-directions. CrB4 demonstrates anisotropic thermal conductivity in all three directions. Following the approach in relevant studies [74, 75], we employed MD simulations to validate the $\kappa_{\rm L}$ values of these conductor candidates. For Be2C, calculations are required for only one direction due to its cubic symmetry. For CN2, we considered only $\kappa_{\rm L}$ along the $z$-axis. In the case of CrB4, the overall $\kappa_{\rm L}$ is determined as the average of the values computed along the three principal directions.
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim$30, $\sim$80, $\sim$150 W/m/K). As shown in Figure 8(e), samples with $\kappa_{\rm L}>100$ W/m/K account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for regions with scarce samples typically exhibit considerable uncertainty [76, 77].
KRbAu2 has the space group symbol $\textit{Fm}\overline{3}\textit{m}$, belonging to the cubic crystal system. For such isotropic thermal insulator candidates, it is generally more efficient and accurate to calculate $\kappa_{\rm L}$ using DFT [78, 79]. As shown in Figure 8(d), the DFT-calculated $\kappa_{\rm L}$ for this structure at 300K is 0.201 W/m/K, while the KAN-predicted value is 0.2169, correctly identifying it as a thermal insulator. In summary, the methods of $k$-means clustering and t-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model. Furthermore, the $\kappa_{\rm L}$ values predicted by KAN exhibit excellent agreement with both MD and DFT validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability.
II.4 Tracing high/low LTC from the perspective of theoretical calculation
Among the validated candidates, considering the isotropy of Be2C and KRbAu2, we take these two as examples to explore the mechanisms influencing high and low $\kappa_{\rm L}$ based on DFT. Their crystal structures are shown in Figures 9(a) and (b), both belonging to the cubic system but exhibiting distinctly different phonon transport behaviors. Be2C adopts a [BeC4] tetrahedral configuration, where each C atom is coordinated with four Be atoms, with a bond length of approximately 1.64 ${\rm\AA}$. These [BeC4] tetrahedral units share edges, forming a continuous three-dimensional covalent network that imparts high rigidity and thermal stability to the crystal. In contrast, KRbAu2 consists of K, Rb, and Au, forming a metal-ion mixed-bond network. The Au atoms exist as [Au2] dimers, where each Au atom is connected via metallic bonds to form a framework structure. K and Rb act as cations filling the interstitial spaces, providing electrostatic stabilization. Due to the large atomic radii of K and Rb, the introduced steric hindrance effect increases lattice flexibility [80].
Figure 9: Based on DFT-based phonon thermal transport analysis. Crystal structures and the projected 2D electron localization function (ELF) diagram of (a) CrB4 and (b) KRbAu2. (c) is the -pCOHP & -IpCOHP for Be-C, K-Au and Rb-Au bonds. In (d) and (e), the left panels depict the phonon dispersion curves of CrB4 and KRbAu2, while the right panels present the corresponding atom-projected PDOS.
The effects of lattice rigidity and flexibility on $\kappa_{\rm L}$ primarily manifest in their influence on phonon group velocity $v_{g}$ and anharmonicity [81]. In Be2C, the strong covalent bonding between Be and C forms a highly rigid lattice, leading to smaller atomic displacements and lower lattice vibrational anharmonicity, thereby reducing three-phonon scattering (primarily suppressing Umklapp scattering) [82]. Additionally, such crystals typically exhibit high elastic moduli, resulting in steep phonon dispersion curves, which lead to higher phonon velocities and enhanced thermal conductivity [83]. In contrast, the softer bonds between K, Rb, and Au make the lattice more deformable. A more flexible lattice experiences larger thermal vibrations, inducing stronger three-phonon and four-phonon scattering, which significantly reduces $\kappa_{\rm L}$. Atomic mass also indirectly influences $\kappa_{\rm L}$ by affecting sound velocity. Be2C consists of light atoms, whereas KRbAu2 is composed entirely of heavy atoms. Specifically, heavy atoms tend to reduce group velocity, thereby exhibiting thermal insulation characteristics, whereas light atoms have the opposite effect [84, 85, 86].
Figure 10: Modal quantities of CrB4 and KRbAu2. (a-c) illustrate the frequency dependence of the average thermal conductivity $\bar{\kappa}_{\lambda}$, group velocity $|\mathbf{v}_{\lambda}|$, and phonon lifetime $\tau_{\lambda}$ for CrB4, respectively; (d-f) present the counterparts for KRbAu2. The scatters are colored according to the band index, with two distinct colormaps used to differentiate the two materials.
To further investigate the impact of chemical bonding on the $\kappa_{\rm L}$ of Be2C and KRbAu2, we employ the Negative Projected Crystal Orbital Hamilton Population (-pCOHP) to visualize their local bonding characteristics. In -pCOHP, positive values indicate bonding states, negative values represent antibonding states, and a value of zero signifies non-bonding interactions. As shown in Figure 9(c), the Fermi level $E_{\rm F}$ lies at 0 eV. For the Be-C bond, bonding states dominate below $E_{\rm F}$ down to -8.7 eV. Strong bonding implies greater lattice rigidity and a steeper potential energy surface [87], leading to higher optical branch vibration frequencies, which suppress phonon scattering and enhance thermal conductivity. This behavior is similar to classical high-$\kappa_{\rm L}$ materials such as diamond and SiC. In KRbAu2, the bonding characteristics differ between the two types of bonds: K-Au is bonding, while Rb-Au is antibonding. The antibonding nature of the Rb-Au bond softens the lattice, enhances phonon scattering, and reduces $\kappa_{\rm L}$. Figure 8(d) confirms the ultra-low $\kappa_{\rm L}$ of KRbAu2, suggesting that its phonon transport properties are primarily dictated by the Rb-Au bond. The -IpCOHP analysis further evaluates global bond strength. Despite K-Au being a bonding interaction, its bond strength remains weak, and the presence of an antibonding Rb-Au bond exacerbates phonon anharmonicity. Weak chemical bonds are often associated with strong phonon anharmonicity [88, 89], which is a key factor underlying the ultra-low $\kappa_{\rm L}$ of KRbAu2.
Figures 9(d) and (e) provide insights into the characteristics of low-energy optical branches and acoustic modes, supporting the previous hypotheses. In the phonon dispersion spectrum of Be2C, all phonon modes exhibit real frequencies, indicating dynamical stability at zero temperature with no tendency for structural collapse. The acoustic branches in the low-frequency region (0 $\sim$ 10 THz) show relatively high group velocities, suggesting potentially high phonon thermal conductivity. In contrast, the optical branches are primarily distributed in the high-frequency range (15 $\sim$ 30 THz) and exhibit a certain degree of flatness, indicating lower group velocities and minimal contribution to heat transport. The Projected Density of State (PDOS) analysis on the right side of Figure 9(d) further reveals that heat-carrying phonons are predominantly below 15 THz, mainly contributed by the heavier Be atoms, while the high-frequency optical branches involve both C and Be atoms. For KRbAu2, the phonon dispersion curve shows that all modes remain within the real frequency range, with no imaginary branches, confirming its dynamical stability at zero temperature. The acoustic branches are concentrated in the low-frequency region (0 $\sim$ 1 THz) and exhibit flat dispersions, resulting in low group velocities. The optical branches are distributed between 1 and 3.5 THz, with some modes displaying significant dispersion. According to the PDOS, the dominant vibrating atoms shift sequentially from Au to Rb to K as frequency increases, consistent with the mass distribution of these elements.
Figure 10(b) and (e) show the frequency dependence of the group velocity $\mathbf{v}_{g}$ at the center of the Brillouin zone for Be2C and KRbAu2, respectively. The group velocity of Be2C exceeds 10 km/s, while that of KRbAu2 does not exceed 1.1 km/s. This significant difference in group velocity can be attributed to differences in bonding and strength, which is consistent with the inference made from the steepness of their phonon spectra. Figure 10(c) and (f) describe the differences in phonon lifetimes between the two materials. In the low-frequency range below 10 THz, Be2C exhibits longer phonon lifetimes, with the longest exceeding 10 ps and the shortest being above 1 ps. In contrast, in KRbAu2, the phonon lifetime decreases sharply with increasing frequency, with $\tau_{\lambda}$ approaching 0.1 ps at $\upsilon_{\lambda}=2{\rm THz}$. It is noteworthy that for certain specific low-frequency modes, the phonon lifetime $\tau_{\lambda}$ of KRbAu2 exceeds that of Be2C, which may be due to special scattering suppression effects, e.g., weaker three-phonon scattering [90]. However, overall, Be2C has a longer and more uniform phonon lifetime, which is favorable for long-range heat transport, while the shorter phonon lifetime of KRbAu2 suppresses $\kappa_{\rm L}$.
III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the $\kappa_{\rm L}$ of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based $\kappa_{\rm L}$ modeling has achieved remarkable accuracy, but improvements in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of $\kappa_{\rm L}$ modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
Our work employed KAN to model $\log(\kappa_{\rm L})$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects and dependencies. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features of $\kappa_{\rm L}$, highlighting its superior interpretability in this context. We combined KAN with CGCNN to construct a two-stage prediction framework, where CGCNN predicts key features, and the KAN pretraining model maps these features to $\log(\kappa_{\rm L})$; for physical properties like $C_{\text{v}}$, which CGCNN struggles to predict accurately, we used DFT calculations to complement the framework. In the case of unlabeled new samples, we applied $k$-means clustering to identify 10 potential thermal conductors/insulators. Among them, the $\kappa_{\rm L}$ values of CN2, CrB4, Be2C, and KRbAu2 were verified by DFT and MD, yielding results highly consistent with deep learning predictions, which sufficiently demonstrate KAN’s excellent extrapolation capability. We visualized the phonon spectra, PDOS, -pCOHP, and other information of Be2C and KRbAu2 using DFT, analyzing the phonon transport mechanisms of both materials from the perspectives of chemical bonding, atomic mass, and structural stability, providing theoretical insights as reference.
IV METHODS
IV.1 Kolmogorov-Arnold Networks (KANs)
The Kolmogorov-Arnold representation theorem [91] provides that any multivariate continuous function defined on a bounded domain can be expressed as a finite composition of continuous functions of a single variable, combined with the operation of addition [92]. For a differentiable function $f:{{\left[0,1\right]}^{n}}\to\mathbb{R}$:
$f\left(\mathbf{x}\right)=f\left({{x}_{1}},...,{{x}_{n}}\right)=\sum\limits_{q=1}^{2n+1}{{{\Phi}_{q}}\left(\sum\limits_{p=1}^{n}{{{\phi}_{q,p}}\left({{x}_{p}}\right)}\right)}$ (7)
where ${{\phi}_{q,p}}:\left[0,1\right]\to\mathbb{R}$ as well as ${{\Phi}_{q}}:\mathbb{R}\to\mathbb{R}$. In this context, $p$ denotes the number of top operators, $q$ denotes the number of bottom operators, and $n$ denotes the number of nodes in the bottom network (which corresponds to the dimensionality of the input features in the input layer), $\mathbf{x}=\left({{x}_{1}},...,{{x}_{n}}\right)$ represents feature vector.
In the original K-A representation theorem, the number of nonlinear layers is limited to 2, and the number of hidden layer nodes is set to $12n+$ 1, which fixes the network structure to $\left[n,2n+1,1\right]$. However, Liu et al. [30] are not constrained by these limitations. In KANs, both the number of layers and the width of the network are arbitrary, which enhances the feasibility of the K-A representation theorem for ML applications. Therefore, for a KAN with ${{n}_{0}}$-D inputs, L layers, and ${{n}_{\text{L}}}=1$-D output, a more precise definition is:
$\text{KAN}\left(\mathbf{x}\right)=\sum\limits_{{{i}_{L-1}}=1}^{{{n}_{L-1}}}{{{\phi}_{L-1,{{i}_{L}},{{i}_{L-1}}}}\left(\sum\limits_{{{i}_{L-2}}=1}^{{{n}_{L-2}}}{\cdots\left(\sum\limits_{{{i}_{2}}=1}^{{{n}_{2}}}{{{\phi}_{2,{{i}_{3}},{{i}_{2}}}}\left(\sum\limits_{{{i}_{1}}=1}^{{{n}_{1}}}{{{\phi}_{1,{{i}_{2}},{{i}_{1}}}}\sum\limits_{{{i}_{0}}=1}^{{{n}_{0}}}{{{\phi}_{0,{{i}_{1}},{{i}_{0}}}}\left({{x}_{{{i}_{0}}}}\right)}}\right)}\right)\cdots}\right)}$ (8)
the expression can be simplified to:
$\text{KAN}\left(\mathbf{x}\right)=\left({{\mathbf{\Phi}}_{L-1}}\circ{{\mathbf{\Phi}}_{L-2}}\circ\cdots\circ{{\mathbf{\Phi}}_{1}}\circ{{\mathbf{\Phi}}_{0}}\right)\mathbf{x}$ (9)
in the expression, ${{\mathbf{\Phi}}_{i}}=\left\{{{\phi}_{q,p}}\right\}$ denotes the i-th layer of the KAN, defined as a tensor of 1-D activation functions ${{\phi}_{q,p}}$. In the implementation details, $\phi\left(x\right)$ is represented as a linear combination of the basis function $b\left(x\right)$ and B-spline function:
$\phi\left(x\right)$ $=\omega\left[b\left(x\right)+\text{spline}\left(x\right)\right]$
$=\omega\left[\frac{x}{1+{{\text{e}}^{-x}}}+\sum\limits_{i}{{{c}_{i}}{{B}_{i}}\left(x\right)}\right]$ (10)
where B-spline function is defined as:
$B_{i,0}(x)$ =$=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1}\\ 0,&\text{otherwise}\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x)$ (11)
In KANs, the coefficients $c$, the degree $k$, and the grid $G$ of the spline functions are all learnable parameters, which are updated through backpropagation.
IV.2 Sure Independence Screening & Sparsifying Operator (SISSO)
SISSO constructs a descriptor vector ${{\mathbf{d}}_{n}}$ using features and operators, and models the target vector $\mathbf{P}$ through a linear combination of $n$-dimensional descriptors. The vector of primary features, ${{\mathbf{\Phi}}_{0}}$, serves as the starting point for constructing descriptors. The operator set is defined as:
${{\hat{H}}^{\left(\text{m}\right)}}\equiv\left\{I,+,-,\times,\div,\exp,\log,\text{abs},\sqrt{{}}{{,}^{-1}}{{,}^{-2}}{{,}^{-3}}\right\}\left[{{\phi}_{1}},{{\phi}_{2}}\right]$ (12)
where ${{\phi}_{1}}$ as well as ${{\phi}_{2}}$ are terms in ${{\mathbf{\Phi}}_{i}}$. The superscript ${}^{\left(\text{m}\right)}$ indicates that SISSO retains only descriptors with physical meaning. For example, features that are added or subtracted must have the same dimensions, and features involved in logarithmic or square root operations cannot be negative. The new features constructed during the $i$-th iteration can be expressed as
${{\Phi}_{i}}=\bigcup\limits_{k}{{{{\hat{h}}}_{k}}^{\left(\text{m}\right)}\left[{{{\hat{\phi}}}_{i}},{{{\hat{\phi}}}_{j}}\right]},\forall{{\hat{h}}_{k}}^{\left(\text{m}\right)}\in{{\hat{H}}^{\left(\text{m}\right)}}\text{ and }\forall{{\phi}_{i}},{{\phi}_{j}}\in{{\Phi}_{i-1}}$ (13)
where ${{\hat{h}}_{k}}^{\left(\text{m}\right)}$ represents single operator of ${{\hat{H}}^{\left(\text{m}\right)}}$, ${{\phi}_{i}}$ and ${{\phi}_{j}}$ are different elements from ${{\Phi}_{i-1}}$. SIS [93] constructs the feature subspace with the highest correlation to the target $\mathbf{P}$ through vector inner products, while SO selects the top $n$ features with the highest relevance using regularization techniques to form the descriptor matrix ${{\mathbf{d}}_{n}}$. The linear coefficients can be approximately determined by solving the equation ${{\mathbf{c}}_{n}}={{\left(\mathbf{d}_{n}^{\text{T}}{{\mathbf{d}}_{n}}\right)}^{-1}}\mathbf{d}_{n}^{\text{T}}\mathbf{P}$, so the model constructed by SISSO can be represented as:
$\mathbf{\hat{P}}={{\mathbf{d}}_{n}}{{\mathbf{c}}_{n}}={{c}_{0}}+{{c}_{1}}{{d}_{1}}+\cdots+{{c}_{n}}{{d}_{n}}$ (14)
where $\mathbf{\hat{P}}$ represents the estimated target vector of the model.
IV.3 Sobol Indices
Sobol is a method for feature importance analysis based on analysis of variance (ANOVA) [60]. It allocates a portion of the total variance to each input variable or its interactions with other variables, thereby providing valuable information about the importance of each input variable [65]. The first-order and total Sobol indices can be defined as follows:
${{S}_{i}}=\frac{\text{Var}\left[{{\text{E}}_{\sim i}}\left[Y|{{X}_{i}}\right]\right]}{\text{Var}\left[Y\right]}$ (15)
$S_{i}^{\text{T}}=\frac{{{\text{E}}_{\sim i}}\left[\text{Var}\left[Y|{{\mathbf{X}}_{\sim i}}\right]\right]}{\text{Var}\left[Y\right]}$ (16)
for a specific value of ${{X}_{i}}$, the value of $Y$ can be determined by averaging the model evaluations over a sample of ${{\mathbf{X}}_{\sim i}}$ while keeping ${{X}_{i}}=x_{i}^{*}$ fixed, where ${{\mathbf{X}}_{\sim i}}$ represents all variables except ${{X}_{i}}$.
IV.4 Kucherenko Indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies [62, 65, 94]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas [95] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure [96], as detailed in Table S2. This approach approximates the correlated variables as independent before calculating the Sobol indices. The Gaussian copula is constructed based on the standard normal distribution, and its mathematical expression is:
$C\left({{u_{1}},{u_{2}},\ldots,{u_{d}};\Sigma}\right)={\Phi_{\Sigma}}\left[{{\Phi^{-1}}\left({{u_{1}}}\right),{\Phi^{-1}}\left({{u_{2}}}\right)\ldots,{\Phi^{-1}}\left({{u_{d}}}\right)}\right]$ (17)
where $u_{i}=F_{i}(x_{i})$ is the Cumulative Distribution Function (CDF) of the marginal distribution, $\Phi^{-1}$ is the inverse of the CDF of the standard normal distribution $\mathcal{N}(0,1)$, $\Phi_{\Sigma}$ is the CDF of a multivariate normal distribution with mean zero and covariance matrix $\Sigma$, and $\Sigma$ is the correlation matrix, describing the dependence structure between the random variables.
IV.5 Dataset construction & feature preprocessing
The dataset used in this work is sourced from the aflowlib.org database [97], comprising a total of 5,578 entries. For feature selection, the focus was primarily on characteristics related to vibrational, thermodynamic, and mechanical properties at room temperature (300K), which are theoretically associated with $\kappa_{\rm L}$ according to first-principles researches [98, 99]. For more details, please refer to Table 2. In the modeling aimed at $\kappa_{\rm L}$, the model predicts the logarithm of $\kappa$ ($\log(\kappa_{\rm L})$) rather than the $\kappa_{\rm L}$ itself. This is because the logarithmic transformation compresses the target value space, thereby enhancing the performance of the ML models [100]. For deep models such as MLPs and KANs, normalization of features is essential as these models rely on gradient descent for parameter optimization. Normalization enhances numerical stability by ensuring that features with different ranges contribute equally to the gradient computations; otherwise, features with larger values might dominate the optimization process [101]. However, this step is unnecessary for XGBoost [102], as it is based on decision trees. Additionally, for interpretability reasons, feature normalization is often omitted in recent works based on SISSO [26, 28, 103]. In this work, we employed Min-Max Normalization, which is expressed as:
$x^{\prime}=\frac{{x-{x_{\min}}}}{{{x_{\max}}-{x_{\min}}}}$ (18)
where $x$ is the primary feature value, ${{x}_{\min}}$ and ${{x}_{\max}}$ are the minimum and maximum values of the feature, respectively, and ${x}^{\prime}$ is the normalized feature value.
IV.6 Implementation of DFT & GPUMD methods
DFT and AIMD calculations are performed using VASP [104, 105]. The Projector-Augmented Wave (PAW) method [106] and the Perdew-Burke-Ernzerhof (PBE) functional [107] within the Generalized Gradient Approximation (GGA) are employed for electron exchange-correlation. The cutoff energy is set to 500 eV, and the electronic convergence threshold is $10^{-8}$ eV.
For the calculation of $\kappa_{\rm L}$, both the phonon BTE method and MD simulations are used. For the stable structures, $\kappa_{\rm L}$ are computed using phono3py [108, 109, 110]. While $\kappa_{\rm L}$ of unstable structures are determined through GPUMD [111] AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. Uniformly spaced sampling is performed, with 250 frames extracted to train the NEP model. The HNEMD method is employed in GPUMD simulations. A Fe value of $1\times 10^{-4}\,{\rm\AA}^{-1}$ has been tested and found to be optimal for this study. Specific heat and phonon dispersions are calculated using Phonopy [112]. The pCOHP analysis is performed using the LOBSTER code [113]. Structural visualizations and ELF analysis are conducted using the VESTA software [114].
IV.7 Data Availability
The original LTC data and the selected features can be downloaded from https://aflowlib.org/ or accessed via the GitHub homepage at https://github.com/FlorianTseng/LTC-modeling.
IV.8 Code Availability
The codes supporting our research are available at https://github.com/FlorianTseng/LTC-modeling.
IV.9 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements.
We would like to acknowledge the National Key Research and Development Program of China (Grant No. 2023YFB4603800), the financial support from the project of the National Natural Science Foundation of China (Grants No.12404062, No.12474093, No.12302220, No.12374091), Translational Medicine and Interdisciplinary Research Joint Fund of Zhongnan Hospital of Wuhan University (Grant NO.ZNJC202235), and the Fundamental Research Funds for the Central Universities (Grant No.2042023kf0109). The numerical calculations in this work have been done on the supercomputing system in the Supercomputing Center of Wuhan University.